OpenStudy (anonymous):

susan is six years older than beth. in 4 years susan will be twice as old as beth how old is susan now?

OpenStudy (anonymous):

let the present age of Beth be x and hence present age of Susan is x+6

OpenStudy (anonymous):

after 4yrs age of Beth is x+4 and that of Susan is (x+6)+4 =x+10 given x+10 = 2(x+4) can you solve for x now..

OpenStudy (anonymous):

What he said^ :)

OpenStudy (anonymous):

Thanks